# Agile Decision-Making and Safety-Critical Motion Planning for Emergency Autonomous Vehicles

### v3

@@ Abstract
Efficiency is critical for autonomous vehicles (AVs), especially for emergency AVs. However, most existing methods focus on regular vehicles, overlooking the distinct strategies required by emergency vehicles to address the challenge of maximizing efficiency while ensuring safety. In this paper, we propose an Integrated Agile Decision-Making with Active and Safety-Critical Motion Planning System (IDEAM). IDEAM focuses on enabling emergency AVs, such as ambulances, to actively attain efficiency in dense traffic scenarios with safety in mind. Firstly, the speed-centric decision-making algorithm named the long short-term spatio-temporal graph-centric decision-making (LSGM) is given. LSGM comprises conditional depth-first search (C-DFS) for multiple paths generation as well as methods for speed gains and risk evaluation for path selection, which presents a robust algorithm for high efficiency and safety consideration. Secondly, with an output path from LSGM, the motion planner reconsiders environmental conditions to decide constraints states for the final planning stage, among which the lane-probing state is designed for actively attaining spatial and speed advantage. Thirdly, under the Frenet-based model predictive control (MPC) framework with final constraints state and selected path, the safety-critical motion planner employs decoupled discrete control barrier functions (DCBFs) and linearized discrete-time high-order control barrier functions (DHOCBFs) to model the constraints associated with different driving behaviors, making the optimal optimization problem convex. Finally, we extensively validate our system using scenarios from a randomly synthetic dataset, demonstrating its capability to achieve speed benefits and assure safety simultaneously.
The simulation video is available at: https://www.youtube.com/watch?v=873BZoQSf-Q
@@ I Introduction
With the advancement of perception and communication technologies, autonomous vehicles (AVs) are increasingly integrating into urban and highway traffic environments. In such highly interactive scenarios, it is anticipated that AVs have the potential to enhance driving safety [1, 2, 3] and efficiency [4, 5, 6], which are particularly critical for emergency AVs such as ambulances, fire trucks, and emergency service auto taxis  [7]. Existing studies predominantly focus on decision-making [8, 9, 10] or motion planning [11, 12, 13], respectively, and others propose integrated systems [14, 15] designed for regular AVs in urban or highway environments. However, the applications of AVs assigned to special tasks like emergency vehicles are rarely paid attention to. Therefore, it is crucial to design a system that integrates decision-making and motion planning to maximize efficiency gains while maintaining safety.
@@
Figure 1: The proposed IDEAM framework. The LSGM decision-making layer is utilized to generate the graph path with the highest efficiency and the final desired vehicle group, while the motion planner actively and safely navigates AV to the desired vehicle group.
Decision-making is a challenging task in highly interactive traffic scenarios where autonomous vehicles (AVs) and human-driven vehicles (HDVs) coexist. Recent research has been dedicated to developing learning-based models for lane changes that take driving styles into account. For instance, in [16], a personalized support vector machine (SVM) is employed to facilitate lane-changing decisions in two-lane scenarios. Additionally, [17] introduced a model based on convolutional neural networks (CNNs) for lane-changing decisions that integrate considerations of driving styles. Such models require a large amount of pre-collected data and computing resources [18]. However, collecting datasets of emergency vehicles in real traffic scenarios is challenging, because emergency vehicles operate in high-risk scenarios such as high-speed driving, and overtaking, making frequent real-world testing unsafe and difficult. Moreover, reinforcement learning (RL)-based approaches are commonly used for decision-making in lane-changing scenarios. For example, deep Q-learning (DQN) serves as a high-level behavior planner for lane changes [19, 20]. However, it is important to note that RL decision makers may pose safety risks due to limitations in exploration, generalization, and the lack of explicit safety constraints. On the other hand, the rule-based models can become increasingly complex and difficult to maintain as the number of vehicles grows. Building upon the method introduced by [18], [7] introduced a decision-making framework with a finite state machine (FSM) aiming for speed benefits. However, this framework’s hand-crafted rules are complex, lack scalability, and may not consistently deliver improvements in speed gains.
Motion planning is a critical part of autonomous driving, responsible for generating a reliable trajectory for AVs based on the outputs from the decision-making layer. A large amount of recent research has been focused on model predictive control (MPC) [21, 22, 23, 24, 25], which is a powerful tool in numerical optimization methods that allows for the application of motion planning. Ensuring vehicle safety is a critical concern in motion planning, typically addressed through constraint-handling methods such as barrier Lyapunov functions (BLFs) [26, 27] and control barrier functions (CBFs) [28, 29, 30, 31]. While MPC optimizes a trajectory considering future states within a defined prediction horizon, CBFs serve as effective safety filters that enforce safety through forward invariance, ensuring that vehicles consistently remain within a designated forward safe set. [32] is the first to combine MPC with discrete control barrier function (DCBF) to ensure safety-critical motion planning. However, this integration leads to a nonlinear programming problem (NLP), which not only increases computational demand but also heightens the risk of settling for local optima or encountering infeasible solutions. To address this issue, [33] introduced discrete-time high-order control barrier function (DHOCBF), a novel and promising approach that renders the MPC-DHOCBF method convex. However, its application is currently limited to scenarios involving only static obstacles.
@@
To address the above-mentioned issues, we propose an Integrated Agile Decision-Making with Active and Safety-Critical Motion Planning System (IDEAM). As shown in Fig. 1, our proposed system consists of a decision-making layer and a motion planning layer. The decision-making is generated by the LSGM algorithm, which is an agile and efficiency-oriented method with strong scalability. The LSGM algorithm solves a graph-based path-finding problem by maximizing speed gains while ensuring safety. Using the C-DFS strategy, the algorithm efficiently identifies paths within the graph, leveraging both short-term and long-term speed benefits. Additionally, by incorporating spatio-temporal risk assessment and gap magnitude judgment, the LSGM algorithm ensures the safety of the selected paths. Ultimately, it provides an optimized and safe path to the motion planning layer. For the motion planning layer, we integrate various constraint states within the MPC framework to formulate different optimization problems, among which the lane-probing constraint state enables AVs to actively explore spatial advantages rather than merely adopting a passive safety mode in response to the environment. Moreover, corresponding to different constraint states, we apply distinct decoupled DCBF and dynamically linearized DHOCBF under the framework, making the optimization problems under these states convex. In summary, the main contributions of this paper are as follows:
@@
  • We introduce a safety-critical motion planner capable of dynamically adjusting constraints based on varying constraint states and actively exploring spatial advantages. The entire motion planning process is framed as convex numerical optimization problems.
@@
The rest of this paper is organized as follows. Section II presents the preliminary of this paper. Section III introduces the Long short-term spatio-temporal graph-centric decision-making algorithm. The motion planning framework is presented in Section IV. In Section V, simulation results of the proposed method are discussed. SectionVI concludes the paper.
II Preliminary
@@
In this section, the preliminary related to the motion planning module is introduced. We present the vehicle dynamics model in Section II-A, while some background of the control barrier function is introduced in Section 3.
II-A Vehicle Dynamics Model
@@
where $v_{x}$ and $v_{y}$ denote the longitudinal velocity and lateral velocity respectively. $w$ represents the yaw rate of the vehicle. $a_{x}$ is the longitudinal acceleration and $\delta_{f}$ denotes steer angle. $F_{fy}$ and $F_{ry}$ are the lateral forces on the front and rear tires while $m$ stands for the mass. $I_{z}$ indicates the moment of inertia. $l_{f}$ and $l_{r}$ refer to the distances from the center of the vehicle to the front and rear axles, respectively.
The Pacejka’s Magic Formula [34] is adopted to model the lateral forces of the tires as follows
$F_{fy}=2D_{f}\sin{({C_{f}\arctan{(B_{f}\alpha_{f})}})},$ (7)
$F_{ry}=2D_{r}\sin{({C_{r}\arctan{(B_{r}\alpha_{r})}})},$ (8)
@@
$\alpha_{f}=\delta_{f}-\arctan{(\frac{l_{f}w+v_{y}}{v_{x}})},$ (9)
$\alpha_{r}=\arctan{(\frac{l_{r}w+v_{y}}{v_{x}})},$ (10)
Figure 3: Visualization of linearized DHOCBF. The linearized safe set, shown in blue, guarantees collision avoidance.
@@
$\mathbf{x}_{k+1}=\mathbf{A}\mathbf{x}_{k}+\mathbf{B}\mathbf{u}_{k}+\mathbf{C},$ (11)
@@
$\mathbf{x}=\begin{bmatrix}v_{x},v_{y},w,s,e_{y},e_{\psi}\end{bmatrix},\mathbf{u}=\begin{bmatrix}a_{x}\delta\end{bmatrix}.$ (12)
II-B Control Barrier Functions
@@
$\dot{\mathbf{x}}=f(\mathbf{x})+g(\mathbf{x})\mathbf{u},$ (13)
@@
When referring to the discrete-time domain, discrete CBF (DCBF) formulation (14) can be shown as follows
$\Delta h(\mathbf{x}_{k},\mathbf{u}_{k})\geq-\gamma h(\mathbf{x}_{k}),0<\gamma\leq 1,$ (15)
@@
In the context of our study, the relative degree [29] of the continuously differentiable function $h:\mathbb{R}^{n}\to\mathbb{R}$ is higher than 1 with respect to the system (13). Consequently, we introduce the discrete-time high-order control barrier function (DHOCBF) approach to address this characteristic.
@@
$C_{m}(t)$ :=$:=\{\mathbf{x}\in\mathbb{R}^{n}:\psi_{m-1}(\mathbf{x},t)\geq 0\},$
@@
$\begin{split}\psi_{i-1}(\mathbf{x}_{t,k})+\sum_{\nu=1}^{i}Z_{\nu,i}(1-\gamma_{i})^{k}\psi_{0}(\mathbf{x}_{t,\nu})&\geq\\ \omega_{t,k,i}Z_{0,i}(1-\gamma_{i})^{k}\psi_{0}(\mathbf{x}_{t,0}),\end{split}$ (18)
@@
III Long short-term spatio-temporal graphcentric decision-making
This section introduces the LSGM algorithm. With the information of nearby vehicles, the LSGM can make lane-changing decisions for the ego vehicle.
III-A Problem Formulation
@@
Figure 4: An illustration of risk assessment within the decision-making time domain in C-DFS algorithm.
III-B Conditional Depth-First Search
@@
Risk Assessment. When searching for connected paths, the safety of connectivity between nodes is crucial for successful and accident-free lane changes. Risk assessment is employed to determine the safety of the connection between adjacent nodes within a path. If this connection is deemed unsafe, the path is subsequently filtered out. As depicted in Fig 4, ensuring a safe connection between two nodes necessitates a sufficient temporal gap $d_{risk}$ between the lead vehicle of the pointing node and the rear vehicle of the pointed node, thus enabling the ego vehicle to pass through securely. During the risk assessment process, when evaluating the connection between each pair of nodes, distinct time domains in the prediction horizon are considered, each characterized by a fixed length. Earlier connections correspond to preceding time domains, while subsequent connections correspond to succeeding time domains. The mathematical formulation of $d_{risk}$ is defined as follows
$d_{risk}^{z,k}=\begin{cases}2l_{diag}+\epsilon&v_{l,\text{ego}}^{z,k}>v_{f,i}^{z,k},\\ 2l_{diag}+n\Delta v+\epsilon&v_{f,i}^{z,k}>v_{l,\text{ego}}^{z,k},\end{cases}$ (20)
@@
  • $\mathbf{LevelCheck}$: verify if a node’s level is below the highest recorded level. Should it be lower, the search for paths passing through this node will be disregarded, an example is presented in Fig 5. This approach ensures that once a path progresses through the spatially advantageous group $2$, it does not revert to group $1$.
@@
Figure 5: An illustration of LevelCheck in C-DFS. LevelCheck prevents paths from reverting from group $2$ to group $1$ by filtering out nodes with lower levels. In this figure, paths starting from $L_{1}$ and ending at $L_{2}$ will not revert back to $R_{1}$ or $C_{1}$ once they progress through $R_{2}$ or $C_{2}$.
@@
III-C Long Short-term Efficiency Group Selector
Much works focus on improving efficiency by following a faster leader vehicle. However, considering only the speed of the leader vehicle is insufficient, as larger spaces in other vehicle groups may provide opportunities for overtaking and other maneuvers beneficial for speed gains. Unlike these methods, we refine assessment by evaluating the predicted spatial positions of surrounding vehicles. The longitudinal position of the leader vehicle in future time frames not only implies information about the vehicle’s speed but also indicates the availability of sufficient space for maneuvers such as overtaking, accelerating, and lane-changing, all aimed at enhancing efficiency. Inspired by this notion, we categorize the efficiency evaluation of a path into two aspects: long-term efficiency and short-term efficiency.

### v4

@@ Abstract
Efficiency is critical for autonomous vehicles (AVs), especially emergency AVs. However, most existing methods focus on regular vehicles, overlooking the different strategies required by emergency vehicles to address the challenge of maximizing efficiency while ensuring safety. In this paper, we propose an Integrated Agile Decision-Making with Active and Safety-Critical Motion Planning System (IDEAM). IDEAM focuses on enabling emergency AVs, such as ambulances, to actively achieve efficiency in dense traffic scenarios with safety in mind. Firstly, the speed-centric decision-making algorithm named the long short-term spatio-temporal graph-centric decision-making (LSGM) is given. LSGM comprises conditional depth-first search (C-DFS) for multiple path generation as well as methods for speed gains and risk evaluation for path selection, which presents a robust algorithm for high efficiency and safety consideration. Secondly, with an output path from LSGM, the motion planner reconsiders environmental conditions to decide constraint states for the final planning stage, among which the lane-probing state is designed for actively attaining spatial and speed advantage. Thirdly, under the Frenet-based model predictive control (MPC) framework with final constraints state and selected path, the safety-critical motion planner employs decoupled discrete control barrier functions (DCBFs) and linearized discrete-time high-order control barrier functions (DHOCBFs) to model the constraints associated with different driving behaviors, making the optimization problem convex. Finally, we extensively validate our system using scenarios from a randomly synthetic dataset, which reveal that IDEAM improves average route progress by approximately $5.25\%$ to $12.93\%$ and increases average speed by about $4.5\%$ to $9.8\%$ compared to the benchmark method, demonstrating its capability to achieve speed benefits and assure safety simultaneously.
Simulation video is available at: https://www.youtube.com/watch?v=873BZoQSf-Q
Our implementation code is available at https://github.com/YimingShu-teay/IDEAM.git.
@@ I Introduction
With the advancement of perception and communication technologies, autonomous vehicles (AVs) are increasingly integrating into urban and highway traffic environments. In such highly interactive scenarios, it is anticipated that AVs have the potential to enhance driving safety [1, 2] and efficiency [3, 4, 5], which are particularly critical for emergency AVs (EAVs). EAVs refer to vehicles designed to handle urgent, time-sensitive tasks such as ambulances, fire trucks, and emergency service auto taxis  [6]. Existing studies predominantly focus on decision-making [7, 8, 9] or motion planning [10, 11, 12], respectively, and others propose integrated systems [13, 14] designed for regular AVs in urban or highway environments, considering efficiency and safety. However, the applications of AVs assigned to special tasks like emergency vehicles are rarely explored. Therefore, it is crucial to design a system that integrates decision-making and motion planning to maximize efficiency gains while maintaining safety. In this paper, we address these gaps by developing an integrated system specifically tailored for emergency AVs. Our work advances in decision-making and motion planning, providing specialized solutions that prioritize high-speed performance.
@@
Figure 1: The proposed IDEAM framework. The LSGM decision-making layer is utilized to generate the graph path with the highest efficiency and the final desired vehicle group, while the motion planner actively and safely navigates the AV to the desired vehicle group.
Decision-making and motion planning are critical for autonomous vehicles, especially in dynamic traffic environments. Learning-based methods for decision-making, such as SVMs [15], CNNs [16], have been explored for lane-changing decisions. However, deploying such methods in EAVs may present challenges, as specialized datasets for priority-driven navigation are scarce, and their collection may pose difficulties. On the other hand, rule-based approaches, for example, FSM frameworks [6], offer interpretability but often struggle with scalability. For motion planning, optimization-based approaches using model predictive control (MPC) [17, 18, 19, 20] augmented with control barrier functions [21] have emerged as a promising safety-critical solution in recent years. However, in dynamic settings, the implicit feasible region in nonlinear programming problems may lead to the risk of local optima [22]. These challenges are exacerbated for emergency vehicles requiring rapid, efficient, and safety-critical responses.
@@
To address the above-mentioned issues, we propose an Integrated Agile Decision-Making with Active and Safety-Critical Motion Planning System (IDEAM). As shown in Fig. 1, our proposed system consists of a decision-making layer and a motion-planning layer. The decision-making is generated by the LSGM algorithm, which is an agile and efficiency-oriented method with strong scalability. The LSGM algorithm solves a graph-based path-finding problem by maximizing speed gains while ensuring safety. Using the C-DFS strategy, the algorithm efficiently identifies paths within the graph, leveraging both short-term and long-term speed benefits. Additionally, by incorporating spatio-temporal risk assessment and gap magnitude judgment, the LSGM algorithm ensures the safety of the selected paths. Ultimately, it provides an optimized and safe path to the motion planning layer. For the motion planning layer, we integrate various constraint states within the MPC framework to formulate different optimization problems, among which the lane-probing constraint state enables AVs to actively explore spatial advantages rather than merely adopting a passive safety mode in response to the environment. Moreover, corresponding to different constraint states, we apply distinct decoupled DCBF and dynamically linearized DHOCBF under the framework, making the optimization problems under these states convex. In summary, the main contributions of this paper are as follows:
@@
  • We introduce a safety-critical motion planner capable of dynamically adjusting constraints based on varying constraint states and actively exploring spatial advantages. The entire motion planning process is framed as a convex numerical optimization problem.
@@
The rest of this paper is organized as follows. Section II reviews relevant work on decision-making and motion planning. Section III presents the preliminary of this paper. Section IV introduces the LSGM algorithm. The motion planning framework is detailed in Section V. Section VI discusses the simulation results of the proposed method. Section VII analyzes the system’s limitations and potential directions for future work. Section VIII concludes the paper.
II Related Works
Decision Making Recently, various decision-making methods have been developed for autonomous vehicles to handle complex interactive scenarios [23]. Traditional traffic models include Minimizing Overall Braking Induced by Lane changes (MOBIL) [24, 25], which improves overall traffic efficiency by minimizing braking impacts during lane changes. However, such models lack the emphasis on pursuing speed benefits, which is essential for EAVs. Rule-based methods, such as Finite State Machine (FSM) [26], [6], explicitly define vehicle behaviors through predefined states and transitions, but their complexity rapidly increases as the number of interacting agents grows, leading to large and complex state tables with limited flexibility. Additionally, Partially Observable Markov Decision Processes (POMDP) [27, 28] and game-theoretic methods [29, 30], are regarded as theoretically well-founded decision-making methods considering interaction and environmental uncertainties among vehicles. However, as the number of interacting vehicles increases, the state or action space may grow combinatorially, raising concerns about computational complexity. To address this, recent works [31, 32], [33] have incorporated learning-based trajectory prediction [34], [35], [36], [37] to replace the behavior simulations. The incorporation of data-driven components into decision-making relies on large-scale datasets. However, most existing datasets are tailored to common urban civilian driving scenarios, limiting the generalizability of such methods to more specialized or high-priority contexts. This challenge similarly affects learning-based decision-making approaches, such as those utilizing CNNs [16] and SVMs [38], [15], [39]. In contrast, reinforcement learning (RL)-based methods [40], [41, 42] enable policy optimization through interaction with simulated environments without relying on labeled datasets, but it typically consider safety only through reward functions, which may be insufficient to address collision risks during exploration.
Motion Planning Applications in lane keeping [43, 44], lane changing [45, 46], and merging [30] have become key research directions in motion planning for autonomous vehicles in recent years. Among various approaches, control barrier functions (CBFs) have emerged as a promising choice, offering safety guarantees via forward invariance to keep the system within predefined safe sets. A CLF-CBF-QP framework was first proposed by [47] for adaptive cruise control, and was later extended by [48], who integrated it with a rule-based switching strategy to enable safe autonomous lane changes. However, the method provides only one-step safety guarantees without considering long-term planning horizons. Model predictive control (MPC), on the other hand, optimizes over a receding horizon, enabling robust trajectory planning with explicit handling of constraints. [22] first introduced the integration of MPC with discrete-time control barrier functions (MPC-DCBF), where the feasibility of the resulting nonlinear programming problem (NLP) was analyzed, highlighting the tradeoff between safety and feasibility. Further applications of MPC-DCBF include [49], which demonstrated its effectiveness in low-speed car racing scenarios. Recently, the linearized discrete high-order control barrier function (DHOCBF) has emerged as an effective solution to the nonlinear programming (NLP) problem, offering fast computational performance while ensuring safety. While it has been applied to static obstacle avoidance, it holds promise for extension to dynamic interactive scenarios.
III Preliminary
@@
In this section, the preliminary related to the motion planning module is introduced. We present the vehicle dynamics model in Section III-A, while some background of the control barrier function is introduced in Section 3.
III-A Vehicle Dynamics Model
@@
where $v_{x}$ and $v_{y}$ denote the longitudinal velocity and lateral velocity, respectively. $w$ represents the yaw rate of the vehicle. $a_{x}$ is the longitudinal acceleration and $\delta_{f}$ denotes steer angle. $F_{fy}$ and $F_{ry}$ are the lateral forces on the front and rear tires, while $m$ stands for the mass. $I_{z}$ indicates the moment of inertia. $l_{f}$ and $l_{r}$ refer to the distances from the center of the vehicle to the front and rear axles, respectively.
Pacejka’s Magic Formula [50] is adopted to model the lateral forces of the tires as follows
$F_{fy}=2D_{f}\sin{({C_{f}\arctan{(B_{f}\alpha_{f})}})}$ (7)
$F_{ry}=2D_{r}\sin{({C_{r}\arctan{(B_{r}\alpha_{r})}})}$ (8)
@@
$\alpha_{f}=\delta_{f}-\arctan{(\frac{l_{f}w+v_{y}}{v_{x}})}$ (9)
$\alpha_{r}=\arctan{(\frac{l_{r}w+v_{y}}{v_{x}})}$ (10)
@@
$\mathbf{x}_{k+1}=\mathbf{A}\mathbf{x}_{k}+\mathbf{B}\mathbf{u}_{k}+\mathbf{C}$ (11)
@@
$\mathbf{x}=\begin{bmatrix}v_{x},v_{y},w,s,e_{y},e_{\psi}\end{bmatrix},\mathbf{u}=\begin{bmatrix}a_{x},\delta\end{bmatrix}$ (12)
III-B Control Barrier Functions
@@
$\dot{\mathbf{x}}=f(\mathbf{x})+g(\mathbf{x})\mathbf{u}$ (13)
@@
When referring to the discrete-time domain, the discrete CBF (DCBF) formulation (14) can be shown as follows
$\Delta h(\mathbf{x}_{k},\mathbf{u}_{k})\geq-\gamma h(\mathbf{x}_{k}),0<\gamma\leq 1$ (15)
@@
In the context of our study, the relative degree [52] of the continuously differentiable function $h:\mathbb{R}^{n}\to\mathbb{R}$ is higher than 1 concerning the system (13). Consequently, we introduce the discrete-time high-order control barrier function (DHOCBF) approach to address this characteristic.
@@
$C_{m}(t)$ :=$:=\{\mathbf{x}\in\mathbb{R}^{n}:\psi_{m-1}(\mathbf{x},t)\geq 0\}$
@@
$\begin{split}\psi_{i-1}(\mathbf{x}_{t,k})+\sum_{\nu=1}^{i}Z_{\nu,i}(1-\gamma_{i})^{k}\psi_{0}(\mathbf{x}_{t,\nu})&\geq\\ \omega_{t,k,i}Z_{0,i}(1-\gamma_{i})^{k}\psi_{0}(\mathbf{x}_{t,0})\end{split}$ (18)
@@
IV Long short-term spatio-temporal graphcentric decision-making
This section introduces the LSGM algorithm. With the information about nearby vehicles, the LSGM can make lane-changing decisions for the ego vehicle.
Figure 3: Visualization of linearized DHOCBF. The linearized safe set, shown in blue, guarantees collision avoidance. The gray area represents the infeasible region segmented by DHOCBF, while the green area denotes the infeasible region delineated by the linearized DHOCBF at time step $k$.
IV-A Problem Formulation
@@
Figure 4: An illustration of risk assessment within the decision-making time domain in the C-DFS algorithm. A node consists of three components: the leading vehicle, the following vehicle, and the gap between them (represented by the gray area in the figure). When C-DFS determines the first node to visit, the time horizon is truncated to the first segment, i.e., $z=1$. Here, $k$ represents the timestep within each time segment $z$. The lower part of the figure illustrates that at $k=2$ in the $z=1$ time segment, the assessment of $d_{\text{risk}}$ is based on the leader vehicle of node $C_{1}$ and the following vehicle of node $R_{1}$.
IV-B Conditional Depth-First Search
@@
Risk Assessment. When searching for connected paths, the safety of connectivity between nodes is crucial for successful and accident-free lane changes. Risk assessment is employed to determine the safety of the connection between adjacent nodes within a path. If this connection is deemed unsafe, the path is subsequently filtered out. As depicted in Fig. 4, ensuring a safe connection between two nodes necessitates a sufficient temporal gap $d_{risk}$ between the lead vehicle of the pointing node and the rear vehicle of the pointed node, thus enabling the ego vehicle to pass through securely. During the risk assessment process, when evaluating the connection between each pair of nodes, distinct time domains in the prediction horizon are considered, each characterized by a fixed length. Earlier connections correspond to preceding time domains, while subsequent connections correspond to succeeding time domains. The mathematical formulation of $d_{risk}$ is defined as follows
$d_{risk}^{z,k}=\begin{cases}2l_{diag}+\epsilon&v_{l,\text{ego}}^{z,k}>v_{f,i}^{z,k}\\ 2l_{diag}+n\Delta v+\epsilon&v_{f,i}^{z,k}>v_{l,\text{ego}}^{z,k}\end{cases}$ (20)
@@
  • $\mathbf{LevelCheck}$: verify if a node’s level is below the highest recorded level. Should it be lower, the search for paths passing through this node will be disregarded, an example is presented in Fig. 5. This approach ensures that once a path progresses through the spatially advantageous group $2$, it does not revert to group $1$.
@@
  • Complexity Analysis: A standard DFS performs an exhaustive search with a worst-case complexity of $O(b^{d})$. In contrast, C-DFS leverages pruning mechanisms, including LevelCheck and RiskAssessment, and benefits from a limited number of candidate nodes per step, which together substantially reduce the search space. Consequently, the practical complexity is $O(\min(P,b^{d},N+E))$, where $P$ is the number of valid paths, $b$ is the branching number, $d$ is the search depth, and $N$ and $E$ denote the number of nodes and edges, respectively. Empirically, C-DFS achieves near-constant time performance with negligible computational cost, supporting its suitability for real-time applications.
Figure 5: n illustration of path generation and LevelCheck in C-DFS. When the start and end nodes are determined, the C-DFS searches for multiple safe paths validated through risk assessment. Taking the path from the start node $L_{1}$ to the end node $L_{2}$ as an example, the figure illustrates multiple possible paths from $L_{1}$ to $L_{2}$. Importantly, LevelCheck prevents paths from reverting from group $2$ to group $1$ by filtering out nodes with lower levels, for example, paths will not revert to $R_{1}$ or $C_{1}$ once they progress through $R_{2}$ or $C_{2}$.
@@
IV-C Long Short-term Efficiency Group Selector
Much works focus on improving efficiency by following a faster leader vehicle. However, considering only the speed of the leader vehicle is insufficient, as larger spaces in other vehicle groups may provide opportunities for overtaking and other maneuvers beneficial for speed gains. Unlike these methods, we refine the assessment by evaluating the predicted spatial positions of surrounding vehicles. The longitudinal position of the leader vehicle in future time frames not only implies information about the vehicle’s speed but also indicates the availability of sufficient space for maneuvers such as overtaking, accelerating, and lane-changing, all aimed at enhancing efficiency. Inspired by this notion, we categorize the efficiency evaluation of a path into two aspects: long-term efficiency and short-term efficiency.